OpenStudy (anonymous):

Antiderivative question ∫(1/√(4-x))dx

OpenStudy (lgbasallote):

\[\int \frac{1}{\sqrt{4-x}}dx\] let \[u = 4 - x\] \[du = -dx\] so the integral becomes \[\implies - \int \frac{1}{\sqrt u} du\] does that help?

OpenStudy (anonymous):

so only use whats inside the root, not the whole denominator. ok. it helps a lot, i was getting a really complicated formula. thank you

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (lgbasallote):

and you can use sqrt (4-x) as u as well although it gets kinda complicated

OpenStudy (lgbasallote):

so for simplicity's sake...it's best to just use the simplest possible substitution